OpenStudy (anonymous):

Hi, I need help on this ? plz.. x - 2y + z = 0 5y - 4z = 7 2z = 4 answer choice: A. x = -4, y = 3, z = 6 B. x = 4, y = -3, z = 2 C. x = 8, y = -3, z = 2 D. x = 4, y = 3, z = 2 help:3

OpenStudy (anonymous):

start at the bottom if \(2a=4\) then what is \(z\) ?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

lol sorry, i meant \(2z=4\) yes \(z=2\)

OpenStudy (anonymous):

Then wat do i do ? :O

OpenStudy (anonymous):

second equation is \[5y-4z=7\] and if \(z=2\) then it is \[5y-4\times 2=7\] or \[5y-8=7\]

OpenStudy (anonymous):

if \(5y-8=7\) then what is \(y\) ?

OpenStudy (anonymous):

How can i find the y ? do i add 8 both sides ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

add 8 to both sides, then divide by 5

OpenStudy (anonymous):

3 ?

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

which pretty much only leaves one answer, but you can find \(x\) using the first equation

OpenStudy (anonymous):

wait, what i got confused ?

OpenStudy (anonymous):

please explain ? (:

OpenStudy (anonymous):

we have \(z=2,y=3\) and your answer choices only have one with these but you can put \[x-2y+z=0 \\x-2\times 3+2=0\\x-6+2=0 \\x-4=0\] and solve for \(x\)

OpenStudy (anonymous):

4 ? @satellite73

OpenStudy (anonymous):

(: